X = X, +*2+1 solve general solution of linear system of NON-hams geneous Diff Equation
Answer #1

te2t x = 2cie 2t - 26e-2t. x = (x) + (x4), IT 7 x₂ = x; -x - 1 x₂ = 2 qe2t - cze 2t - - - ((, e2t + cze 2t - 1 +-&) - 2 * x2